OpenStudy (anonymous):

Find the equation of the ellipse with the following properties. The ellipse with x-intercepts at (4, 0) and (-4, 0), y-intercepts at (0, 9) and (0, -9), and center at (0, 0).

OpenStudy (anonymous):

@TheWeeknd16 @Smita12 @Awolflover1

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Isaiah.Feynman

Nnesha (nnesha):

\[\huge\rm \frac{ (x-h)^2 }{ a^2}+\frac{ (y-k)^2 }{ b^2 }=1\] standard form equation of ellipse so first of all plug in center points (h,k) (0,0)

Nnesha (nnesha):

ello??

OpenStudy (anonymous):

hi :P

OpenStudy (anonymous):

X-( u calll me aunty meri bachi tum to 5 sal ki ho na

Nnesha (nnesha):

*participate* plz it is really important even if i ask what is 1+1 = ?

OpenStudy (anonymous):

mene participate kiya tha chalo byeee

Nnesha (nnesha):

anyways y-intercept is when x=0 x-intercept when y =0 so by looking at the graph you will figure out which number is belongs to x and which one is y-intercept recall: first number in order pair is x and 2nd one is y (x,y)